Solve the formula for h
show all the steps
L = 3πch + 3πc^2

Answer by oscargut(2103)

L = 3πch + 3πc^2
L = 3πc(h +c)
h+c = l/(3πc)
h = l/(3πc) -c
I can help you with more questions at: mthman@gmail.com
Thanks